Psalm 37

Of David

Do not fret because of evil men or be envious of those who do wrong; for like the grass they will soon wither, like green plants they will soon die away.

Trust in the LORD and do good; dwell in the land and enjoy safe pasture. Delight yourself in the LORD and he will give you the desires of your heart.

Commit your way to the LORD; trust in him and he will do this: He will make your righteousness shine like the

dawn, the justice of your cause like the noonday sun.

Be still before the LORD and wait patiently for

him; do not fret when men succeed in their ways, when they carry out their wicked schemes.

Refrain from anger and turn from wrath; do not fret–it leads only to evil. For evil men will be cut off, but those who hope in the LORD will inherit the

land.

A little while, and the wicked will be no more; though you look for them, they will not be found. But the meek will inherit the land and enjoy great peace.

The wicked plot against the righteous and gnash their teeth at them; but the Lord laughs at the wicked, for he knows their day is coming.

The wicked draw the sword and bend the bow to bring down the poor and needy, to slay those whose ways are upright. But their swords will pierce their own hearts, and their bows will be broken.

Better the little that the righteous have than the wealth of many wicked; for the power of the wicked will be broken, but the LORD upholds the righteous.

The days of the blameless are known to the LORD, and their inheritance will endure forever. In times of disaster they will not wither; in days of famine they will enjoy plenty.

But the wicked will perish: The LORD’s enemies will be like the beauty of the

fields, they will vanish–vanish like smoke.

The wicked borrow and do not repay, but the righteous give generously; those the LORD blesses will inherit the land, but those he curses will be cut off.

If the LORD delights in a man’s way, he makes his steps firm; though he stumble, he will not fall, for the LORD upholds him with his hand.

I was young and now I am old, yet I have never seen the righteous forsaken or their children begging bread. They are always generous and lend freely; their children will be blessed.

Turn from evil and do good; then you will dwell in the land forever. For the LORD loves the just and will not forsake his faithful ones.

They will be protected forever, but the offspring of the wicked will be cut off; the righteous will inherit the land and dwell in it forever.

The mouth of the righteous man utters wisdom, and his tongue speaks what is just. The law of his God is in his heart; his feet do not slip.

The wicked lie in wait for the righteous, seeking their very lives; but the LORD will not leave them in their power or let them be condemned when brought to trial.

Wait for the LORD and keep his way. He will exalt you to inherit the land; when the wicked are cut off, you will see it.

I have seen a wicked and ruthless man flourishing like a green tree in its native soil, but he soon passed away and was no more; though I looked for him, he could not be found.

Consider the blameless, observe the upright; there is a future for the man of peace. But all sinners will be destroyed; the future of the wicked will be cut off.

The salvation of the righteous comes from the

LORD; he is their stronghold in time of trouble. The LORD helps them and delivers them; he delivers them from the wicked and saves them, because they take refuge in him.